MICO Meeting 20 - 28th April 2008
- RAL: A. Nicholls, C. Rogers, M.Ellis (MOM), A. Blondel, T. Bradshaw, K. Long, C. Macwaters, T. Hayler
- Phone: K. Tilley, M. Zisman, S. Griffiths
Action Items (A Nichols)
Target - Still waiting for a reply from the magnet vendors.
Decay magnet - cooldown is proceeding well.
South wall infrastructure is mainly here and being installed, need three weeks to finish.
False floor design is stalled due to need for force modelling to be solved.
Need to know what to do very soon. Tom will work on this.
North wall order is about to be placed, have permission to do so without a tender.
Badly mixed concrete is on a flat part of the floor and will be a simple digging out and re-pouring operation. The ramp is unaffected.
The solenoid racks are shown on the drawings but need better labeling, that will be done.
A. Nichols has spoken to Maurizio and the help from Hammamatsu has not completely solved the problem, still likely to be delayed until middle of May.
Maurizio will determine the size of the box(es) required for the signal cables to come to RAL early.
The cables have BNC connectors for signal and SHV for the high voltage.
ISIS shutdown dates revised.
Useful target meeting this morning.
Target coil thermal test is still not done.
Target flanges and tubes are at RAL and have been given to the Sheffield people.
Some issues with the new target shafts but not thought to be serious.
Permanent magnets quotes are in process.
Coils will arrive tomorrow.
Cooling test with the decay solenoid, the projected timescale is to be around He temperatures by Wednesday and have power to the solenoid by the end
of this week with the end of the test by the end of next week.
False floor design is held up for the magnetic modelling, Tom will progress that.
False floor installation date has passed.
North wall magnetic shield support order is about to be placed.
PPS - nothing to report
Controls - nothing to report
DAQ - still need to interpret document from JSG
Spectrometer solenoid - nothing new to report
Focus coil order - requestition has been signed, contract sent to Tesla.
Detectors - Alan is on furlough. Alain suggested a replica of GVA2 to help determine efficiencies.
Possibly we could borrow the D1 and D2
detectors before the tracker test.
ISIS issues (C. Rogers)
New ISIS dates are now firm and will be posted.
Long, productive meeting this morning.
Now have permission to run a few thousand pulses with the current target in the current user run.
Have permission to open the gate valve once this is signed off by Paul Wright.
Need to operate the new target for a few days prior to installation in ISIS.
ISIS require a signal from MICE for the target pulse and target position.
They will provide the signal from the beam loss monitors.
Machine activation measurements - ISIS suggested to place samples of different materials near the target to study activation as well as suggestions for how to make a generic activation measurement.
As we get more pulses this will be understood.
Perhaps include target magnets in the activation test in ISIS.
Targets (C Booth)
Already covered in Milestones.
Cryogenic systems (M. Courthold)
Testing the Linde plant and decay solenoid at the moment.
After this test, will accept the plant and will change the compressor.
Liquid H2 systems are coming along.
Still have a bottleneck due to resources going to decay solenoid and Linda plant issues.
Will be back onto that once this test is complete.
Wing has produced a note.
Infrastructure ( T Hayler)
Contractrors that are doing the south magnetic wall are doing the work today.
Will take approximately 3 weeks to put the verticval struts up before we can mount the steel.
Steel order is timed to turn up in time for this.
Should run seamlessly for the next few months.
False floor order is more likely to be placed in July and have it ready to be installed in September/October to have the north and south walls up first.
This affects step 3, so we may have to buy and install the first part earlier.
Cages, DSA roof and escape routes are set for the May sutdown.
Infrastructure (S Griffiths)
Currently looking at the overall cable management for the tracker, CKOV, TOFs and KL and in general the requirements for the diagnostics.
This has come out to be a
large number of cables and if all are going back to the LCR then we have 500 cables.
A quick calculation shows that we can fit 200 cables per hole (based on the TOF cables).
There is a hole for HT and another for personnel safety, another for diagnostics, etc.
Initially there was only one for diagnostics but we will need at least two now.
One of these holes will possibly become available once the HT system is moved out of the control room later on.
K.L. pointed out we can remove the wooden door and take some cables through the labyrinth.
S.G. has a cable inventory that is being built up.
When that is done, then each cable can be identified with its diameter, type and connector to be put on a drawing to be connected with the holes in the wall.
Majority of people want HV connections in the control room, but KL is being supplied with 16m length cables that will not reach the control room, so the HV will have to be in the MICE hall.
On the KL system, have a cable tray from the cable runs on the wall right across the MICE hall and will pass straight through the metal shielding wall.
Don't want to install this tray and then have to remove it for the wall, so it looks like there is a clash.
Can use the floor duct to come up near Q9. Need to check with Ludovico if this is OK.
Not ruling out bringing the cable tray though the metal wall.
PPS (J Alexander)
No progress to report until the shutdown.
Cable access routes to the DSA have been approved.
Controls (B Martlew)
Spectrometer Solenoid (S Virostek)
Slides on the web presented by MZ.
RFCC module - plan for a review in the CM at DL in June. (currently planned for 4th June 2pm-6pm).
KL asked about the bolt holes on the upstream plate of the upstream solenoid.
Wing was to contact Steve Virostek to sort this out.
Not clear that these will be put in advance.
MZ will mention this to SV.
SG asked about the supply voltage - they should operate on 220 or 440 V.
AFC Module (W Lau)
Slides on the web.
Detectors (A Bross)
A Bross is on furlough this week.
Maurizio - visit with Hamamatsu: showed them the problem and they have taken notes.
Apparantly the probvlems are mainly connected to the active divider.
We are the first experiment using this type of PMT with this design of the active divider.
What has been delivered was a "beta" version.
Agreed that on the new delivery there will be 50 tubes to be delivered.
Will try to assemble TOF0 with the PMTs that will be delivered at the beginning of May and the best of the old PMTs.
In the meantime the mounting collars have been changed and this has helped a little.
Plan to use the May shutdown to deliver TOF0, but cannot give a precise date just yet.
AN requested one week notice for the shipment. MB agreed to give at least one week notice.
KL suggested making a stand for TOF0 that would go behind the concrete in case we can't get it inside the DSA.
Details on the web.
ME will request safety and operational documents for MICE to be collected on the web.